OpenStudy (anonymous):

If Brian is a 46% free throw shooter, what is the probability of him making at least 3 free throws out of 6?

OpenStudy (anonymous):

.46^3 + .54^3 *.46^3 + .54^2 * .46^3 + .54 * .46^3 = About 19.4%? Is that right?

OpenStudy (anonymous):

Can you show me a different way to do it if I'm right?

OpenStudy (anonymous):

at least 3 is annoying 3 or 4 or 5 or 6 or you could compute 0 or 1 or 2 and subtact the number from 1 your choice

OpenStudy (anonymous):

you are forgetting to multiply by the number of ways to do it

OpenStudy (anonymous):

\[\dbinom{n}{k}p^k(1-p)^{n-k}\] is what you need, and it looks like what you have except for the \[\dbinom{n}{k}\] part

OpenStudy (wasiqss):

yup satellite is rite, .46^3+.46^4+.46^5+.46^6

OpenStudy (anonymous):

all 6 is \[.46^6\] 5 out of six is \[5\times .46^5\times .54\]

OpenStudy (anonymous):

why 5 times that?

OpenStudy (anonymous):

exactly 4 is \[\dbinom{6}{4}\times .46^4\times .54^2\]

OpenStudy (anonymous):

oh i get it 6 choose 5 = 5

OpenStudy (anonymous):

six shots, you want five there are bernoulli trials yes six choose five

OpenStudy (anonymous):

finally exactly 3 is \[\dbinom{6}{3}\times.46^3\times .54^3\] you good from there right? calculate and add

OpenStudy (anonymous):

thanks satellite, so it was my equation except i have to multiply by the number of ways to do each term in that?

OpenStudy (anonymous):

well it also looks like your exponents were a little messed up

OpenStudy (anonymous):

they should add to six for one thing

OpenStudy (anonymous):

okay, what i was thinking was that after missing 2 then making 3 it didnt matter what the next one was.

OpenStudy (anonymous):

say i want exactly 2 out of 6 i think \[.46^2\] two successes \[.54^4\] 4 failures \[\dbinom{6}{2}\] number of ways to get 2 successes out of 6 so answer would be \[\dbinom{6}{2}\times .46^2\times .54^4\]

OpenStudy (anonymous):

oh but it does matter, because if you are computing exactly 3 successes, another success would make it 4!

OpenStudy (anonymous):

but it said at least 3

OpenStudy (anonymous):

at least 3 means 3 or 4 or 5 or 6 compute each separately

OpenStudy (anonymous):

okay, thanks satellite

OpenStudy (anonymous):

yw